Graph the following equation, label any intercepts. y = ln(x - 1) + 2
Answer by fractalier(6550)

This is the standard y = ln x graph, shifted one point to the right and two points up.
To find out where it crosses the x-axis, set y = 0 and solve for x:
0 = ln(x - 1) + 2
-2 = ln(x - 1)
e^(-2) = x - 1
x = 1 + e^(-2)
The graph never touches the y-axis since x can never be zero here...